OpenStudy (anonymous):

Solve 3x^2 + 21x = -30

OpenStudy (lilg132):

do you know the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (lilg132):

ok let me write it out

OpenStudy (lilg132):

actually here this is the formula take a look at it http://2.bp.blogspot.com/_V8KsSIiGjBk/SsACMEj73KI/AAAAAAAAFIU/vNtErLdchMw/s1600/Quadratic+Formula.gif

OpenStudy (lilg132):

now in order to use this formula you need to make the equation you have = 0 do you know how to do that?

OpenStudy (immanuelv):

its \[x=(-21\pm \sqrt{21^{2}-(4(3)(30))} )/2(3)\]

OpenStudy (lilg132):

if you don know how to make your equation = 0 i can help you

OpenStudy (anonymous):

these are the answer options x = 6 and x = 5 x = –5 and x = –2 x = 7 and x = –5 x = –6 and x = –2

OpenStudy (lilg132):

ok but first you have to make the equation 3x^2 + 21x = -30 = 0 instead of = -30 do you get how to do that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

could you explain please?

OpenStudy (lilg132):

ok so the equation 3x^2 + 21x = -30 at the moment = -30 to get rid of -30 and make it = 0 we add 30 to the other side which will make the equation 3x^2 + 21x + 30 = 0

OpenStudy (anonymous):

alright.. then what do i do next?

OpenStudy (lilg132):

ok so now to use the formula we need to know the values of a, b and c to input that in the formula

OpenStudy (anonymous):

k...

OpenStudy (lilg132):

and remember the equation should be in this order: ax^2 + bx + c = 0

OpenStudy (anonymous):

okay

OpenStudy (lilg132):

which it already is

OpenStudy (lilg132):

the equation is 3x^2 + 21x + 30 = 0 and the value of A is the first number we have there which is 3 the value of B is the second number we have there which is 21 the value of C is the third number we have there which is 30 remember if any of the numbers were - it would be a minus value e.g. if the third number was -30, then C would be -30 so now we have the values a = 3 b = 21 c = 30

OpenStudy (lilg132):

now we need to place this in the formula do you understand how to do that if not tell me i will go through the steps with you

OpenStudy (anonymous):

no i dont know how to find that

OpenStudy (lilg132):

ok have you taken a look at the formula i showed you on http://2.bp.blogspot.com/_V8KsSIiGjBk/SsACMEj73KI/AAAAAAAAFIU/vNtErLdchMw/s1600/Quadratic+Formula.gif

OpenStudy (anonymous):

okay got it

OpenStudy (lilg132):

ok so first thing we need to do is but the value for b where it says -b so we know the value for b is 21 but because it says -b we have to change 21 to -21 (but remember if the value of b was -21 and we had to put it where -b is it would change to a positive) so do you understand that so far that where it says -b we put -21

OpenStudy (anonymous):

yes i understand

OpenStudy (lilg132):

ok so now where the square root sign is for b^2 we put in 21^2 or 441

OpenStudy (lilg132):

and for 4ac we put in 4 * a * c which is 4 * 3 * 30

OpenStudy (lilg132):

so now we have the values in the square root sign we can work out what should be there by doing b^2 - 4ac which is 441 - 360 which would be \[\sqrt{81}\]

OpenStudy (lilg132):

do you understand everything so far?

OpenStudy (anonymous):

yes

OpenStudy (lilg132):

ok so square root of 81 is 9 so we got rid of the square root sign and we have just 9 there so our formula is now \[x -21 \pm 9 \over 2a\]

OpenStudy (lilg132):

the value for 2a is simply 2 * a which is 2 * 3 which is 6

OpenStudy (lilg132):

so our formula now looks like this \[x = {-21 \pm 9 \over 6}\]

OpenStudy (lilg132):

now you would have noticed there is a + and - sign in between -21 and 9 which means we have to do both things and this will give us two different answers

OpenStudy (lilg132):

do you know how to work this out now?

OpenStudy (anonymous):

yes thanks

OpenStudy (lilg132):

dont forget to click good answer :) and your welcome